OpenStudy (anonymous):

what are the inflection points f(x)=e^(-x^2)

OpenStudy (anonymous):

Find its second derivative and make it equal to zero and solve for x

zepdrix (zepdrix):

Look back at your previous question ames, those were your inflection points. :) The x='s.

OpenStudy (anonymous):

i could have sworn we just did this one

OpenStudy (anonymous):

thats what i thought but then how do i tell if it concaved up or concaved down?

OpenStudy (anonymous):

second derivative is \[f''(x)=(4x^2-2)e^{-x^2}\] which is zero if \[4x^2-2=0\] or \[x=\pm\frac{1}{\sqrt2}\]

OpenStudy (anonymous):

concave up if \(f''(x)>0\)

OpenStudy (anonymous):

ignore the \(e^{-x^2}\) part as it is always posive solve \(4x^2-2>0\)

OpenStudy (anonymous):

\[y=4x^2-2\] is a parabola that faces up, so it is negative between the zeros and positive outside of them

OpenStudy (anonymous):

i.e. \(f''(x)<0\) on the interval \((-\frac{1}{\sqrt2}, \frac{1}{\sqrt2})\)

OpenStudy (anonymous):

positive otherwise this more or less clear?

OpenStudy (anonymous):

yes thank you.. i got a bit confused